# PyKirigami: An interactive Python simulator for kirigami metamaterials

### v2

@@ Abstract
In recent years, the concept of kirigami has been used in creating deployable structures for various scientific and technological applications. While the design of kirigami metamaterials has been widely studied, the simulation of the deployment and shape transformation process is less explored. In this work, we develop PyKirigami, an efficient Python-based open-source computational tool for the deployment simulation of kirigami metamaterials. In particular, our tool is capable of simulating both two- and three-dimensional deployments of a large variety of kirigami metamaterials with different geometric, topological, and physical properties. The tool is also highly interactive and customizable, allowing the user to perform deployment simulations of kirigami structures under different environmental setups and material properties. Altogether, our work paves a new way for the modelling and design of shape-morphing mechanical metamaterials.
@@ 1 Introduction
Kirigami, the traditional paper-cutting art, has recently become a subject of great interest in science and engineering [1, 2, 3, 4, 5, 6]. In particular, a wide range of deployable structures with different shape-morphing effects have been developed by introducing cuts on a sheet of materials and connecting different components using suitable joints. Because of their great flexibility and shape-shifting property, kirigami-based metamaterials have been widely applied in soft electronics [7], energy storage [8], and robotics [9].
@@
In this work, we develop PyKirigami, an open-source Python-based simulator, for simulating the 2D and 3D deployment of general kirigami metamaterials (Fig. 1). Specifically, our simulator utilizes the PyBullet physics engine [40] and is capable of simulating 2D-to-2D, 2D-to-3D, and 3D-to-3D kirigami deployments. It also supports the configurations of environmental effects and stabilization mechanisms for modelling different scenarios. It further allows dynamic control of the boundary conditions throughout the deployment process, thereby providing great flexibility in achieving different desired deployment effects. Using PyKirigami, one can easily simulate the deployment and analyze the properties of a wide range of kirigami structures.
@@ 2.1 Overview
The goal of our PyKirigami simulator is to simulate the deployment process of kirigami metamaterials, typically from a compact initial state to a deployed configuration in either 2D or 3D. More specifically, the functionalities of PyKirigami (Fig. 1a) can be categorized as follows:
@@
The core concepts in PyKirigami are presented in further detail in the following sections. See also the SI Section S1–S2 for more detailed explanations.
@@ 2.2 Geometric Representation: From 2D Polygons to 3D Tiles
More mathematically, let $\{\mathcal{T}_{i}\}_{i=1}^{N}$ be the collection of all kirigami tiles in a kirigami structure, where each tile $\mathcal{T}_{i}$ is initially defined by a sequence of vertices $\{v_{i,1},v_{i,2},\dots,v_{i,n_{i}}\}$, where $v_{i,j}\in\mathbb{R}^{3}$ and $n_{i}$ is the number of vertices for tile $\{\mathcal{T}_{i}\}$ (see also Fig. 2a). To create a more physically realistic model, each planar tile is extruded into a 3D rigid body, or “brick,” with a user-defined thickness, $t$. This process is detailed in Algorithm 1 and illustrated in Fig. 2b.
@@ 2.4 Force Models for Actuation and Dynamics
3:  Retrieve current position, orientation from Bullet engine.
@@ 3 Results
To demonstrate the versatility and effectiveness of PyKirigami, we present a curated set of examples grouped by deployment dimensionality. These cases showcase how the simulator’s automated deployment mechanisms, environmental controls, and interactive features are applied to solve distinct challenges in 2D and 3D kirigami systems. For detailed simulation parameters and per-example configurations, see SI Section S3. See also SI Video S1–S4 for the videos of the deployment process of various 2D and 3D kirigami examples.
@@
For completeness, while all examples in Fig. 3 use automated actuation (radial or target-based forces applied each simulation step), PyKirigami also supports a no-force mode and fully interactive runs in the GUI (manual actuation, pin/unpin, pause/reset). In SI Section S3, we show how to apply these functionalities to achieve the target shape of Fig. 3b.
3.2 Deployment Dynamics: Implicit Hinge Elasticity and Active Control
A primary application of PyKirigami is not only to visualize deployment but also to understand and control the underlying physics. Using the compact reconfigurable square-to-disk structure shown in Fig. 3b as a case study, we demonstrate how the simulator reveals emergent physical properties of the kirigami system. We first identify an implicit elastic behavior that drives its passive dynamics and then show how active control is essential for achieving stable, deterministic reconfiguration (see also SI Section S4 for more information).
3.2.1 Implicit Elasticity and Passive Dynamics
When simulating the square-to-disk structure in a low-dissipation environment (near-zero damping, no friction) without any external actuation from the user’s force model, we observe that it does not remain static. Instead, it spontaneously begins to oscillate between its “square” and “disk” configurations, as shown in Fig. 4.
Figure 4: Passive dynamics driven by implicit elasticity. The plot shows the translational (blue) and rotational (red) kinetic energy of the 25-tile square-to-disk structure over 200 simulation steps ($\Delta t=1/240$ s per step). Five key snapshots are provided in the inset to visualize the kirigami configuration with its energy state: (i) A compact square (high implicit potential, low kinetic). (ii) Expansion driven by conversion of potential to kinetic energy, peaking at a fully open state, where the negative-space parallelograms are nearly perfect squares (angles $\approx\pi/2$). (iii) Collapse into the compact disk state, dissipating some energy via self-collision. (iv) Re-expansion, again fueled by stored potential energy. (v) Return to square configuration.
This behavior is caused by the point-to-point constraint solver, which creates a restoring force implicitly to maintain vertex coincidence at the joints. This mechanism is functionally equivalent to having implicit torsional springs at the hinges. At the start of the simulation (step 0), the structure is initialized in a compact state with zero kinetic energy. This compressed geometry, however, places the torsional springs under high tension, effectively storing a large amount of potential energy. In the very first time step, the compressed springs apply strong corrective impulses that perform work on the tiles, causing a rapid conversion of this stored potential into the large kinetic energy that initiates the deployment, as seen in the sharp jump from step 0 in the plot.
The resulting dynamics, visualized on a linear scale in Fig. 4, are therefore analogous to those of a passive, elastic, and metastable system. This simulation reveals a powerful physical intuition for the structure’s intrinsic energy landscape and the significant potential energy barrier between its stable states. While small-scale kinetic energy fluctuations persist throughout the simulation due to the continuous work of the constraint solver (a phenomenon clearly visible on a logarithmic scale; see SI Section S4), they are orders of magnitude smaller than the primary energy of deployment. For practical control, we find that combining modest passive dissipation (damping and friction) with a deterministic active control scheme is the most effective strategy for overcoming the main energy barrier while suppressing these residual oscillations, as we demonstrate next.
3.2.2 Stabilization via Active Control
For many applications, a deployed structure must reach a stable final state with vanishing kinetic energy. While the implicit elasticity provides a natural deployment mechanism, relying solely on passive dissipation (material damping or friction) is difficult to tune: high dissipation traps the structure before it reaches the target, whereas low dissipation permits oscillations. We therefore employ PyKirigami’s target-based actuation to deterministically reach and stabilize the disk configuration.
To make the comparison across actuation regimes fair and interpretable, we fix a canonical disk target extracted from the passive run at the 75th step in Fig. 4. We then sweep the vertex-target spring stiffness $k_{s}\in\{50,100,1000\}$ with a fixed velocity-proportional damping $\mu_{v}=8.5$ (force_damping). Unless otherwise stated, the environment uses linear and angular damping of 0.1 and ground friction of 0.5. We obtain the following outcomes:
@@ 3.2.2 Stabilization via Active Control
    Underactuated (Low Stiffness, $k_{s}=50$, Fig. 5a): The actuation is weak. The peak kinetic energy imparted to the system remains significantly below the potential energy barrier identified in Fig. 4. As predicted, the structure lacks the energy to make the full transition and becomes trapped in an intermediate state.
@@
    Critically Actuated (Medium Stiffness, $k_{s}=100$, Fig. 5b): The actuation provides just enough energy for the system’s peak kinetic energy to surpass the barrier. The structure successfully reaches the target disk configuration and, with the help of both active and passive damping, settles into a stable final state. This represents an efficient and well-controlled deployment.
@@
Figure 5: Parametric study of active control for stabilizing the square-to-disk deployment. In all cases, the target is the compact disk state, and actuation damping is fixed ($\mu=8.5$). Each panel shows snapshots of the deployment correlated with the system’s kinetic energy over time. aUnderactuated ($k_{s}=50$): Weak spring forces are insufficient to overcome internal and environmental resistance, trapping the structure in an intermediate state. bCritically Actuated ($k_{s}=100$): The actuation is sufficient to achieve the target configuration, where the structure stabilizes after a final self-collision. cOveractuated ($k_{s}=1000$): Strong forces cause a rapid deployment but induce persistent, high-frequency oscillations (chatter) around the target state, preventing full stabilization.
@@ 4 Discussion
Data Availability
The code and data are available on GitHub at https://github.com/andy-qhjiang/PyKirigami.
Supplementary Information
Appendix S1 Overview
PyKirigami is an open-source Python toolbox for simulating the 2D and 3D deployment of general kirigami metamaterials. The simulator utilizes the PyBullet physics engine and is capable of simulating 2D-to-2D, 2D-to-3D, and 3D-to-3D kirigami deployments with interactive controls. The toolbox is freely available on GitHub:
@@
Appendix S2 Implementation Walkthrough
S2.1 Coordinate System Management
@@
Figure S1: Coordinate systems in PyBullet. The fixed global axes $(X_{w},Y_{w},Z_{w})$ define the world coordinate system. Each kirigami tile (represented by a box) has its own local coordinate system $(X_{l},Y_{l},Z_{l})$ attached to its center. A corner vertex of the blue box, such as $P$, has constant local coordinates $\mathbf{p}_{l}=P-O_{1}$ in its body’s frame, but its world coordinates $\mathbf{p}_{w}$ change as the body moves.
See Fig. S1 for an illustration. For convenience, we provide helper functions that implement the local-world transforms:
@@
S2.2 Object Creation in PyBullet
@@
As described in the main text (see also main text Fig. 2), we define the tile’s outward unit normal $\mathbf{n}$ in world coordinates and extrude along $\mathbf{n}$ to obtain the top face, while the original tile becomes the bottom face. In Fig. S2, we further illustrate this idea for 3D kirigami structures. Specifically, this local-normal convention is essential in 3D-to-3D deployments, where the world $z$-axis generally does not align with the tile’s normal.
Figure S2: An illustration of face extrusion for 3D kirigami structures. Here, note that the quadrilateral patches on opposite faces of the cube are congruent, but their orientations are different. As the outward normal flips ($\mathbf{n}\mapsto-\mathbf{n}$), the vertex order must be reversed to maintain outward-facing normals. If the top face uses $(v_{0},v_{1},v_{2},v_{3})$ in counter-clockwise order (viewed from outside), the corresponding bottom face should be indexed $(v_{0},v_{3},v_{2},v_{1})$.
@@
to p.createCollisionShape. PyBullet automatically computes the convex hull of the vertices to form a robust and efficient, though invisible, collision mesh (Fig. S3). Below is an example of the collision shape creation for a triangular tile:
@@
Figure S3: Brick creation in the local frame. a, Extrusion: We compute the world-space top vertices $v_{j}^{t}=v_{j}+t\cdot\mathbf{n}$ and the base origin $O$ (tile center), then convert both bottom and top vertices to local coordinates via $u_{j}=v_{j}-O$ and $u_{j}^{t}=v_{j}^{t}-O$. b, Mesh construction: We build a convex collision vertex cloud (local) and a per-vertex-normal visual mesh (local) with cap and side faces.
@@
The visual shape provides a detailed mesh for rendering sharp edges and correct lighting. The key to achieving this is understanding that PyBullet’s graphics API assigns normals on a per-vertex basis. To create flat-shaded faces with hard edges, we must often duplicate geometric vertices in our data buffers, pairing each copy with the correct face normal. We illustrate this by building a single rectangular side face of the brick illustrated in Fig. S3b, defined by local bottom vertices $(u_{0},u_{1})$ and top vertices $(u_{0}^{t},u_{1}^{t})$. Below is an example:
@@
S2.3 Creating Inter-Tile Constraints

### v3

# PyKirigami: An interactive Python simulator for kirigami structures

@@ Abstract
In recent years, the concept of kirigami has been used in creating deployable structures for various scientific and technological applications. While high-fidelity Finite Element Analysis (FEA) is the standard for analyzing stress distributions and material deformation, it is computationally intensive and often ill-suited for the rapid exploration of vast kinematic configuration spaces. In this work, we develop PyKirigami, a lightweight, open-source Python framework for the efficient deployment simulation of kirigami structures. Unlike continuum mechanics solvers, PyKirigami models tessellations as articulated rigid-body networks, allowing for the real-time simulation of global deployment trajectories and volumetric transformations. The tool incorporates collision detection and interactive actuation, enabling users to validate folding paths and identify geometric locking states in both 2D and 3D topologies. This framework serves as a fast kinematic prototyping tool for kirigami structures, allowing researchers to verify deployment mechanics and self-contacts prior to performing detailed mechanical analysis or physical fabrication.
@@ 1 Introduction
Kirigami, the traditional paper-cutting art, has recently become a subject of great interest in science and engineering [1, 2, 3, 4, 5, 6]. In particular, a wide range of deployable structures with different shape-morphing effects have been developed by introducing cuts on a sheet of materials and connecting different components using suitable joints. Because of their great flexibility and shape-shifting property, kirigami-based structures have been widely applied in soft electronics [7], energy storage [8], and robotics [9].
@@
In this work, we develop PyKirigami, an open-source Python-based simulator, for simulating the 2D and 3D deployment of general kirigami structures (Fig. 1). Specifically, our simulator utilizes the PyBullet physics engine [40] and is capable of simulating 2D-to-2D, 2D-to-3D, and 3D-to-3D kirigami deployments. It also supports the configurations of environmental effects and stabilization mechanisms for modelling different scenarios. It further allows dynamic control of the boundary conditions throughout the deployment process, thereby providing great flexibility in achieving different desired deployment effects. Using PyKirigami, one can easily simulate the deployment and analyze the properties of a wide range of kirigami structures.
@@ 2.1 Overview
The goal of our PyKirigami simulator is to simulate the deployment process of kirigami structures, typically from a compact initial state to a deployed configuration in either 2D or 3D. More specifically, the functionalities of PyKirigami (Fig. 1a) can be categorized as follows:
@@
The core concepts in PyKirigami are presented in further detail in the following sections. See also Appendix A–B for more detailed explanations.
@@ 2.2 Geometric Representation: From 2D Polygons to 3D Tiles
More mathematically, let $\{\mathcal{T}_{i}\}_{i=1}^{N}$ be the collection of all kirigami tiles in a kirigami structure, where each tile $\mathcal{T}_{i}$ is initially defined by a sequence of vertices $\{v_{i,1},v_{i,2},\dots,v_{i,n_{i}}\}$, where $v_{i,j}\in\mathbb{R}^{3}$ and $n_{i}$ is the number of vertices for tile $\{\mathcal{T}_{i}\}$ (see also Fig. 2a). To create a more physically realistic model, each planar tile is extruded into a 3D rigid body, or “brick,” with a user-defined thickness, $t$. This process is detailed in Algorithm 1 and illustrated in Fig. 2b. See also Appendix B for more details of the geometric representation.
@@ 2.4 Force Models for Actuation and Dynamics
3:  Retrieve current position, orientation from the Bullet engine.
@@ 3 Results
To demonstrate the versatility and effectiveness of PyKirigami, we present a curated set of examples grouped by deployment dimensionality. These cases showcase how the simulator’s automated deployment mechanisms, environmental controls, and interactive features are applied to solve distinct challenges in 2D and 3D kirigami systems. For detailed simulation parameters and per-example configurations, see Appendix C. See also Supplementary Video S1–S4 for the videos of the deployment process of various 2D and 3D kirigami examples.
@@
For completeness, while all examples in Fig. 3 use automated actuation (radial or target-based forces applied each simulation step), PyKirigami also supports a no-force mode and fully interactive runs in the GUI (manual actuation, pin/unpin, pause/reset). In Appendix C, we show how to apply these functionalities to achieve the target shape of Fig. 3b.
3.2 Geometric Fidelity of Deployment
PyKirigami utilizes a maximal-coordinate formulation where joint closures are enforced via an impulse-based solver (Projected Gauss-Seidel). Unlike reduced-coordinate approaches that strictly enforce kinematic loops, this iterative method inherently permits constraint violations (drift) between connected vertices. To rigorously assess the impact of this numerical relaxation on kinematic accuracy, we perform a quantitative validation on two distinct kirigami structures: a parametric rectangular tessellation kirigami structure and a complex square-to-disk kirigami structure. In both cases, the simulated deployment trajectories are compared directly against exact analytical solutions derived from the geometric design parameters.
3.2.1 Benchmark Case 1: Rectangular Tessellation
We first consider a standard $M\times N$ rectangular tessellation kirigami structure with deployment angle $\theta$ as a benchmark. The theoretical vertex position $p^{\text{th}}_{i}$ at $\theta$ can be derived easily (see Appendix D). We initialize a $4\times 4$ grid in PyKirigami and actuate it from $\theta=0$ to $\theta=\pi/2$. At each time step, we quantify the drift by computing the Euclidean error $\|p^{\text{sim}}_{i}-p^{\text{th}}_{i}\|$ for all vertices $i$.
Figure 4: Quantitative validation of PyKirigami using the rectangular tessellation kirigami structure. The plot shows the maximal and mean vertex error versus the deployment angle $\theta$. Insets show a comparison of the simulated geometry (gray solid) against the theoretical trajectory (red wireframe) at two points. The left inset shows tessellation at the maximum deviation ($\theta\approx 0.04$) with the magnified view highlighting the drift. The right inset displays the fully deployed state, in which the error is negligible.
As shown in Fig. 4, the error profile demonstrates the high fidelity of the impulse-based solver. The simulation initializes with a transient spike in maximum deviation $2.5\%$ as the solver resolves the ill-conditioned constraints near the compact state. Crucially, the mean relative error remains exceptionally low throughout the entire deployment process. This confirms that the transient spike is isolated to a few specific tiles and joints and does not propagate through the system. The negligible mean error demonstrates that PyKirigami maintains vertex-level coherence, verifying the implementation of the topology constraint.
3.2.2 Benchmark Case 2: Square-to-disk Structure
Having validated the solver’s precision on the above benchmark case, we consider analyzing the deployment of the more complex square-to-disk structure as shown in Fig. 3b. As explained in the theoretical construction of the structure in [25], the deployment of it is also a single-degree-of-freedom mechanism (see Appendix D) where the global radius $R$ is a deterministic function of the tile rotation angle $\theta$. Therefore, in Fig. 5 we plot the global radius against the local deployment angle obtained by Pykirigami and compare the result with the analytical result in [25], from which we can see that the discrete data samples obtained by PyKirigami at different time points of the deployment (red circles) show excellent agreement with the theoretical curve $R(\theta)$ (blue solid line). While micro-deviations exist at the hinge level (drift), the global topology remains locked to the correct deployment path, validating the tool for macroscopic shape analysis.
Figure 5: Quantitative validation of PyKirigami using the square-to-disk kirigami structure. Here, we plot the global radius $R$ versus the deployment angle $\theta$ for both the simulations obtained by PyKirigami (red dots) and the analytical result by [25] (blue solid line). The PyKirigami simulation data align closely with the analytical prediction, confirming that the solver captures the correct kinematic mode. The insets show four snapshots from the simulation.
3.3 Actuation Tuning and Numerical Convergence
As demonstrated in the rectangular tessellation benchmark case (Fig. 4), the solver enforces finite collision margins at the joints, which behave like a repulsive contact force to prevent interpenetration. Consequently, the actuation mechanism must be carefully tuned. If the force is too weak, the structure “jams” prematurely, far from the physical limit. If the force is too strong, it fights the collision engine.
To quantify this operational window, we perform a parametric study of the stiffness $k_{s}$ using the square-to-disk example. We target the fully compact disk state and vary the stiffness $k_{s}\in\{50,100,1000\}$ while maintaining spring damping $\mu_{v}=8.5$, environmental damping of $0.1$ and ground friction of $0.5$. The deployment dynamics are analyzed in Fig. 6.
@@ 3.3 Actuation Tuning and Numerical Convergence
    Underactuated (Low Stiffness, $k_{s}=50$, Fig. 6a): As the structure approaches the compact state, the actuation force by spring is insufficient to overcome the cumulative resistance of ground friction and joint contact impulse. Consequently, the structure jams in an intermediate state, failing to reach the physical packing limit. This highlights the necessity of a minimum actuation threshold to get the desired deployment.
@@
    Critically Actuated (Medium Stiffness, $k_{s}=100$, Fig. 6b): The actuation force is strong enough to compress the tiles against their collision boundaries, effectively closing the gaps to the solver’s limit. Simultaneously, the force is low enough that the system’s damping can dissipate the kinetic energy upon arrival. The structure converges rapidly to the target state and remains stable, demonstrating that PyKirigami can achieve maximal compaction without instability.
@@
Figure 6: Parametric study of active control for stabilizing the square-to-disk deployment. The square-to-disk structure is driven toward the compact disk state under varying spring stiffness $k_{s}$. aUnderactuated ($k_{s}=50$): Weak spring forces are insufficient to overcome internal and environmental resistance, trapping the structure in an intermediate state. bCritically Actuated ($k_{s}=100$): The actuation is sufficient to achieve the target configuration, where the structure stabilizes after a final self-collision. cOveractuated ($k_{s}=1000$): Excessive stiffness causes a rapid deployment but induces persistent, high-frequency oscillations (chatter) around the target state, preventing full stabilization.
@@ 4 Discussion
Data Availability The code and data are available on GitHub at https://github.com/andy-qhjiang/PyKirigami.
Appendix
Appendix A Overview
PyKirigami is an open-source Python toolbox for simulating the 2D and 3D deployment of general kirigami structures. The simulator utilizes the PyBullet physics engine and is capable of simulating 2D-to-2D, 2D-to-3D, and 3D-to-3D kirigami deployments with interactive controls. The toolbox is freely available on GitHub:
@@ Appendix A Overview
Appendix B Implementation Walkthrough
B.1 Coordinate System Management
@@
Figure 7: Coordinate systems in PyBullet. The fixed global axes $(X_{w},Y_{w},Z_{w})$ define the world coordinate system. Each kirigami tile (represented by a box) has its own local coordinate system $(X_{l},Y_{l},Z_{l})$ attached to its center. A corner vertex of the blue box, such as $P$, has constant local coordinates $\mathbf{p}_{l}=P-O_{1}$ in its body’s frame, but its world coordinates $\mathbf{p}_{w}$ change as the body moves.
See Fig. 7 for an illustration. For convenience, we provide helper functions that implement the local-world transforms:
@@
B.2 Object Creation in PyBullet
@@
As described in the main text (see also main text Fig. 2), we define the tile’s outward unit normal $\mathbf{n}$ in world coordinates and extrude along $\mathbf{n}$ to obtain the top face, while the original tile becomes the bottom face. In Fig. 8, we further illustrate this idea for 3D kirigami structures. Specifically, this local-normal convention is essential in 3D-to-3D deployments, where the world $z$-axis generally does not align with the tile’s normal.
Figure 8: An illustration of face extrusion for 3D kirigami structures. Here, note that the quadrilateral patches on opposite faces of the cube are congruent, but their orientations are different. As the outward normal flips ($\mathbf{n}\mapsto-\mathbf{n}$), the vertex order must be reversed to maintain outward-facing normals. If the top face uses $(v_{0},v_{1},v_{2},v_{3})$ in counter-clockwise order (viewed from outside), the corresponding bottom face should be indexed $(v_{0},v_{3},v_{2},v_{1})$.
@@
to p.createCollisionShape. PyBullet automatically computes the convex hull of the vertices to form a robust and efficient, though invisible, collision mesh (Fig. 9). Below is an example of the collision shape creation for a triangular tile:
Figure 9: Brick creation in the local frame. a, Extrusion: We compute the world-space top vertices $v_{j}^{t}=v_{j}+t\cdot\mathbf{n}$ and the base origin $O$ (tile center), then convert both bottom and top vertices to local coordinates via $u_{j}=v_{j}-O$ and $u_{j}^{t}=v_{j}^{t}-O$. b, Mesh construction: We build a convex collision vertex cloud (local) and a per-vertex-normal visual mesh (local) with cap and side faces.
@@
The visual shape provides a detailed mesh for rendering sharp edges and correct lighting. The key to achieving this is understanding that PyBullet’s graphics API assigns normals on a per-vertex basis. To create flat-shaded faces with hard edges, we must often duplicate geometric vertices in our data buffers, pairing each copy with the correct face normal. We illustrate this by building a single rectangular side face of the brick illustrated in Fig. 9b, defined by local bottom vertices $(u_{0},u_{1})$ and top vertices $(u_{0}^{t},u_{1}^{t})$. Below is an example:
@@
B.3 Creating Inter-Tile Constraints